After selling tickets at school, Jeff had $6.00 in dime and quarters. If Jeff had 39 coins in total, how many of each coin did he have?
Can you help me with this?

Answer by stanbon(75887)

After selling tickets at school, Jeff had $6.00 in dime and quarters. If Jeff had 39 coins in total, how many of each coin did he have?
--------------
Quantity Equation: d + q = 39
Value Equation...:10d + 25q = 600 cents
----------------------------------
Modify:
10d + 10q = 10*39
10d + 25q = 600
-----------------------
Subtract 1st from 2nd to get:
15q = 600-390
q = 14 (# of quarters)
----------------
Since d + q = 39, d + 14 = 39
So, d = 15 (# of dimes)
==========================================
Cheers,
Stan H.